Opened 12 years ago

Closed 11 years ago

#49 closed flight processing (fixed)

GB2007/05, flight day 103a/2007, Loch Leven

Reported by: mggr Owned by: mggr
Priority: alpha 4 high Milestone: 2007 data processing completion
Component: Processing: general Keywords: GB2007/05
Cc: Other processors:

Description (last modified by anee)

Data location: ~arsf/arsf_data/in_progress/2007/flight_data/uk/GB2007_05-2007_103a_Loch_Leven

Data arrived from ARSF via initial SATA disk transfer in June.

Scientific details: Data to inform models of potentially toxic cyanobacterial blooms in lakes. See ~arsf/arsf_data/in_progress/2007/ARSF_Applications-GB_2007/GB07-05_*pdf

This is one of three lakes overflown.

Sensors

  • ATM (Delivered 23/April/2008)
  • Eagle (Delivered 19/March/2008)
  • Hawk (Delivered 23/April/2008)

Quicklook

Pixel size is 5

Attachments (4)

103aatm.jpg (926.7 KB) - added by anee 12 years ago.
Pixel size is 5
103aatm_line1.jpg (780.2 KB) - added by anee 12 years ago.
103aatm_line2.jpg (814.6 KB) - added by anee 12 years ago.
103aatm_line3.jpg (799.1 KB) - added by anee 12 years ago.

Change History (36)

comment:1 Changed 12 years ago by mggr

  • Owner changed from amro to anee

comment:2 Changed 12 years ago by anee

all lines processed fine and images generated.
appear to be timing issues on the hawk lines, -0.1 - -0.075 offset appears to be a fairly good fix
cuttoff on images, dem doesnt cover whole area

comment:3 Changed 12 years ago by anee

-fixed dems

-timing issues on the hawk lines

line 1, GPTOFF -1 - -0.75
line 2, GPTOFF ~ -0.125
line 3, possible small ofset needed

-timing issues on eagle lines

line, 2, GTPOFF -0.05
line, 3 possible small ofset needed

comment:4 Changed 12 years ago by anee

tile numbers: no00 no10 no20 nt09 nt19 nt29

comment:5 Changed 12 years ago by anee

-line 3 for both eagle and hawk, GPTOFF ~ -0.1 looks reasonable

-note: atm data still needs to be processed

comment:6 Changed 12 years ago by mggr

Vectors for this area at ~arsf/vectors/from_akw/loch_leven

comment:7 Changed 12 years ago by anee

Atms now processed, should now be a case of selecting exact GPTOFFs for each line.

comment:8 Changed 12 years ago by anee

Ticket is awaiting any improvements to boresight calibration figures for better alignment of the lines (mainly for eagle).

comment:9 Changed 12 years ago by anee

All lines processed for atm, eagle and hawk using appropriate GPTOFF figures and new boresight numbers.

Changed 12 years ago by anee

Pixel size is 5

comment:10 Changed 12 years ago by mggr

  • Description modified (diff)

Added quicklook to description

Changed 12 years ago by anee

Changed 12 years ago by anee

Changed 12 years ago by anee

comment:11 Changed 12 years ago by anee

Processed flight data located in: /users/rsg/anee/scratch_space/GB2007_05-2007_103a_Loch_Leven

Delivery directory created and filled. ATM scripts are current and admin/readme.txt contains info on required GPTOFFs for various lines.

comment:12 Changed 12 years ago by anee

  • Owner changed from anee to mggr

comment:13 Changed 12 years ago by mggr

  • Status changed from new to assigned

Copied into ~arsf/arsf_data/in_progress/2007/flight_data/uk/GB2007_05-2007_103a_Loch_Leven, checked (only minor corrections), looks good :)

Ready to burn to DVD.

comment:14 Changed 12 years ago by mggr

  • Owner changed from mggr to anee
  • Status changed from assigned to new

comment:15 Changed 12 years ago by mggr

  • Description modified (diff)

ATM data delivered (adding note to description)

comment:16 Changed 12 years ago by mggr

  • Description modified (diff)

Oop, delivered 19th, not 20th.

comment:17 Changed 11 years ago by anee

Processed Eagle flight data located in: /users/rsg/anee/scratch_space/GB2007_05-2007_103a_Loch_Leven

Delivery directory created and filled.

comment:18 Changed 11 years ago by anee

  • Owner changed from anee to benj

comment:19 Changed 11 years ago by benj

  • Owner changed from benj to anee

Checked, seems fine. Rsynced to ~arsf/arsf_data/in_progress/2007/flight_data/uk/GB2007_05-2007_103a_Loch_Leven

comment:20 Changed 11 years ago by anee

Eagle data sent to Andrew Tyler on disk 5.

comment:21 Changed 11 years ago by anee

  • Owner changed from anee to benj

Processed Eagle flight data located in: /users/rsg/anee/scratch_space/GB2007_05-2007_103a_Loch_Leven

Delivery directory created (20080403)

comment:22 Changed 11 years ago by anee

ATM data, not Eagle.

comment:23 Changed 11 years ago by benj

  • Owner changed from benj to anee

Checked and rsynced to ~arsf/arsf_data/in_progress/2007/flight_data/uk/GB2007_05-2007_103a_Loch_Leven

comment:24 Changed 11 years ago by anee

  • Owner changed from anee to mark1

Processed Hawk data located in: /users/rsg/anee/scratch_space/GB2007_05-2007_103a_Loch_Leven

Delivery directory created (20080410)

comment:25 Changed 11 years ago by mark1

  • Owner changed from mark1 to anee

Has been checked, Read_me.txt edited for typos. Example commands tested (without DEM) and data looks ok. Copied to ~arsf/arsf_data/in_progress/2007/flight_data/uk/GB2007_05-2007_103a_Loch_Leven/

Ready to deliver.

comment:26 Changed 11 years ago by anee

ATM and Hawk data sent to PI.

comment:27 Changed 11 years ago by anee

  • Description modified (diff)

comment:28 Changed 11 years ago by anee

  • Description modified (diff)

comment:29 Changed 11 years ago by anee

  • Description modified (diff)

comment:30 Changed 11 years ago by mggr

Prepared for archiving (DTG, NEODC), moved to archive area (~arsf/arsf_data/complete-ready_to_archive/).

comment:31 Changed 11 years ago by anee

  • Owner changed from anee to mggr

comment:32 Changed 11 years ago by mggr

  • Resolution set to fixed
  • Status changed from new to closed

Data sent to NEODC for archiving.

Note: See TracTickets for help on using tickets.